### v1

# Aggregating Local and Global Features via Selective State Spaces Model for Efficient Image Deblurring

Abstract
Image deblurring is a process of restoring a high quality image from the corresponding blurred image. Significant progress in this field has been made possible by the emergence of various effective deep learning models, including CNNs and Transformers. However, these methods often face the dilemma between eliminating long-range blur degradation perturbations and maintaining computational efficiency, which hinders their practical application. To address this issue, we propose an efficient image deblurring network that leverages selective structured state spaces model to aggregate enriched and accurate features. Specifically, we design an aggregate local and global block (ALGBlock) to capture and fuse both local invariant properties and non-local information. The ALGBlock consists of two blocks: (1) The local block models local connectivity using simplified channel attention. (2) The global block captures long-range dependency features with linear complexity through selective structured state spaces. Nevertheless, we note that the image details are local features of images, we accentuate the local part for restoration by recalibrating the weight when aggregating the two branches for recovery. Experimental results demonstrate that the proposed method outperforms state-of-the-art approaches on widely used benchmarks, highlighting its superior performance.
Index Terms:
Image deblurring, selective state space, features aggregation
I Introduction
Image deblurring aims to recover a latent sharp image from its corrupted counterpart. Due to the ill-posedness of this inverse problem, many conventional approaches [1, 2] address this by explicitly incorporating various priors or hand-crafted features to constrain the solution space to natural images. Nonetheless, designing such priors proves challenging and lacks generalizability, which are impractical for real-world scenarios.
Stimulated by the success of deep learning for high-level vision tasks, numerous data-driven methods have resorted CNN as a preferable choice and develop kinds of network architectural designs, including encoder-decoder architectures [3, 4, 5], multi-stage networks [6, 7], dual networks [8, 9, 10], generative models [11, 12, 13], and so on. While the convolution operation effectively models local connectivity, its intrinsic characteristics, such as limited local receptive fields and independence of input content, hinder the model’s ability to eliminate long-range dependency features. To alleviate such limitations, various transformer variants [14, 15, 16, 17, 18] have been applied to image deblurring and have achieved better performance than the CNN-based methods as they can better model the non-local information. However, image deblurring often deals with high-resolution images, and the attention mechanism in Transformers incurs quadratic time complexity, resulting in significant computational overhead. To alleviate computational costs, some methods [16, 19] opt to apply self-attention across channels instead of spatial dimensions. However, this approach fails to fully exploit the spatial information, which may affect the deblurring performance. While other methods [18, 15] utilize non-overlapping window-based self-attention for single image deblurring, the coarse splitting approach still falls short in fully exploring the information within each patch.
Figure 1: Computational cost vs. PSNR of models on the GoPro dataset [20]. Our ALGNet achieve the SOTA performance.
Structured state space models [21, 22, 23], notably the enhanced version Mamba, have recently emerged as efficient frameworks due to their ability to capture long-range dependencies with linear complexity. However, Mamba’s [21] recursive image sequence processing method tends to neglect local pixels, while the abundance of hidden states in the state space equation often results in channel redundancy, thereby impeding channel feature learning. Given the critical importance of local and channel features in image deblurring, directly applying the state space model often leads to poor performance.
Taking into account the above analyses, a natural question arises: Is it feasible to design a network that efficiently aggregates local and global features for image deblurring? To achieve this objective, we propose ALGNet, with several key components. Firstly, we develop an efficient global block that utilizes selective structured state spaces to capture long-range dependency information essential for high-quality image deblurring. Secondly, to address the issues of local pixel forgetting and channel redundancy in the state space equation, we design a local block to model local connectivity using simplified channel attention. Thirdly, as image details primarily consist of local features, we emphasize the importance of the local block in restoration by recalibrating the weights through the learnable factor when aggregating the two blocks for recovery. Finally, to alleviate training difficulty, we implement multiple scales of input and output modes. As illustrated in Figure 1, our ALGNet model achieves state-of-the-art performance while preserving computational efficiency compared to existing methods.
The main contributions of this work are:
  1. 1.
    We propose an efficiently aggregates local and global features network (ALGNet) that leverages selective structured state spaces model for image deblurring.
  2. 2.
    We design an aggregate local and global block (ALGBlock) consisting of both global and local components. The global block utilizes selective structured state spaces model to capture long-range dependency information. And the local block models local connectivity using simplified channel attention and address the issues of local pixel forgetting and channel redundancy in the state space equation.
  3. 3.
    We present an aggregate methods that emphasize the importance of the local block in restoration by recalibrating the weights through the learnable factor.
  4. 4.
    Extensive experiments demonstrate that our ALGNet achieves favorably performance against state-of-the-art methods.
II Related Work
II-A Hand-crafted prior-based methods.
Due to the image deblurring ill-posed nature, many conventional approaches [1, 2] tackle this problem by relying on hand-crafted priors to constrain the set of plausible solutions. However, designing such priors is a challenging task and usually lead to complicated optimization problems.
II-B CNN-based methods.
With the rapid advancement of deep learning, instead of manually designing image priors, lots of methods [4, 6, 8, 9, 10, 11, 12, 13] develop kinds of deep CNNs to solve image deblurring. To better explore the balance between spatial details and contextualized information, MPRNet [6] propose a cross-stage feature fusion to explore the features from different stages. MIRNet-V2 [24] introduces a multi-scale architecture to learn enriched features for image restoration. IRNeXt [3] rethink the convolutional network design and exploit an efficient and effective image restoration architecture based on CNNs. NAFNet [4] analyze the baseline modules and presents a simplified baseline network by either removing or replacing nonlinear activation functions. SFNet [25] and FSNet [26] design a multi-branch dynamic selective frequency module and a multi-branch compact selective frequency module to dynamically select the most informative components for image restoration. Although these methods achieve better performance than the hand-crafted prior-based ones, the intrinsic properties of convolutional operations, such as local receptive fields, constrain the models’ capability to efficiently eliminate long-range degradation perturbations.
II-C Transformer-based methods.
Due to the content-dependent global receptive field, the transformer architecture [27] has recently gained much popularity in image restoration [28, 15, 16, 17, 18, 29, 30], demonstrating superior performance compared to previous CNN-based baselines. IPT [28] employs a Transformer-based multi-head multi-tail architecture, proposing a pre-trained model for image restoration tasks. However, image deblurring often deals with high-resolution images, and the attention mechanism in Transformers incurs quadratic time complexity, resulting in significant computational overhead. In order to reduce the computational cost, Uformer [18], SwinIR [30] and U2former [15] computes self-attention based on a window. Nonetheless, the window-based approach still falls short in fully exploring the information within each patch. Restormer [16] and MRLPFNet [19] compute self-attention across channels rather than the spatial dimension, resulting in the linear complexity. However, this approach fails to fully exploit the spatial information. FFTformer [14] explores the property of the frequency domain to estimate the scaled dot-product attention, but need corresponding inverse Fourier transform, leading to additional computation overhead.
II-D State Space Models.
State space models [31, 32, 21, 33, 22, 23] have recently emerged as efficient frameworks due to their ability to capture long-range dependencies with linear complexity. S4 [31] is the first structured SSM to model long-range dependency. S5 [32] propose the diagonal SSM approximation and computed recurrently with the parallel scan. Mega [33] introduced a simplification of S4 [31] to be real- instead of complex- valued, giving it an interpretation of being an exponential moving average. SGConv [34] and LongConv [35] focus on the convolutional representation of S4 and create global or long convolution kernels with different parameterizations. Mamba [21] propose a selective mechanism and hardware-aware parallel algorithm. Many vision tasks start to employ Mamba to tackle image classification [36, 37], image segmentation [38], and so on.
In this work, we leverage selective structured state spaces model to capture long-range dependency information. In order to address the issues of local pixel forgetting and channel redundancy in the state space equation, we model local connectivity using simplified channel attention.
III Method
In this section, we first outline the overall pipeline of our ALGNet. Subsequently, we delve into the details of the proposed ALGBlock which contains global block and local block, and features aggregation (FA).
Figure 2: Overall architecture of ALGNet.
III-A Overall Pipeline
The overall pipeline of our proposed ALGNet, shown in Figure 2, adopts a single U-shaped architecture for image restoration.
Given a degraded image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, ALGNet initially applies a convolution to acquire shallow features $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$ ($H,W,C$ are the feature map height, width, and channel number, respectively). These shallow features undergo a four-scale encoder sub-network, progressively decreasing resolution while expanding channels. It’s essential to note the use of multi-input and multi-output mechanisms for improved training. The low-resolution degraded images are incorporated into the main path through the Convs and concatenation, followed by convolution to adjust channels. The in-depth features then enter a middle block, and the resulting deepest features feed into a four-scale decoder, gradually restoring features to the original size. During this process, the encoder features are concatenated with the decoder features to facilitate the reconstruction. Finally, we refine features to generate residual image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$ to which degraded image is added to obtain the restored image: $\mathbf{\hat{I}}=\mathbf{X}+\mathbf{I}$. It’s important to note that the three low-resolution results are solely used for training. To facilitate the frequency selection process, we optimize the proposed network with the following loss function:
$L=\sum_{i=1}^{4}(L_{char}(\hat{I_{i}},\overline{I}_{i})+\delta L_{edge}(\hat{I_{i}},\overline{I}_{i})+\lambda L_{freq}(\hat{I_{i}},\overline{I}_{i}))$ (1)
where $i$ denotes the index of input/output images at different scales, $\overline{I}_{i}$ denotes the target images and $L_{char}$ is the Charbonnier loss:
$L_{char}=\sqrt{||\hat{I_{i}}-\overline{I}_{i}||^{2}+\epsilon^{2}}$ (2)
with constant $\epsilon$ empirically set to $0.001$ for all the experiments. $L_{edge}$ is the edge loss:
$L_{edge}=\sqrt{||\triangle\hat{I_{i}}-\triangle\overline{I}_{i}||^{2}+\epsilon^{2}}$ (3)
where $\triangle$ represents the Laplacian operator. $L_{freq}$ denotes the frequency domains loss:
$L_{freq}=||\mathcal{F}(\hat{I}_{i})-\mathcal{F}(\overline{I}_{i})||_{1}$ (4)
where $\mathcal{F}$ represents fast Fourier transform, and the parameters $\lambda$ and $\delta$ control the relative importance of loss terms, which are set to $0.1$ and $0.05$, respectively.
III-B Global Block
Long-range dependency information is important for image deblurring, and transformer-based methods [16, 14, 19, 15, 18] are starting to become mainstream due to the limited receptive field of CNNs. However, image deblurring often deals with high-resolution images, and the attention mechanism in Transformers incurs quadratic time complexity, resulting in significant computational overhead. Although some design window-based attention or compute self-attention across channels rather than the spatial dimension, this methods may cause the information loss. To this end, we design a global block with linear complexity to image deblurring, as shown in Figure 2(c). Specifically, given an input feature $F_{in}$, we first reshape and normalization it by layer norm (LN). Then, it will go through two parallel branches. In the first branch, the feature channel is expanded by a linear layer, SiLU activation function, together with the SSM layer. In the second branch, the features channel is also expanded by a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with the element-wise multiplication. Finally, the channel number is projected back and reshape to the original size. The total process can be defined as :
$F_{in}^{1}$ $=SiLU(Linear(Reshape(LN(F_{in}))),$ (5)
$F_{in}^{2}$ $=SSM(SiLU(W_{1D}^{1}(Linear(Reshape(LN(F_{in})))))$
$F_{G}$ $=Reshape(Linear(F_{in}^{1}\otimes F_{in}^{2}))$
III-C Local Block
Although the selective state space model [21] has the ability of global effective receptive field as well as linear computational complexity, it can result in the problem of local pixel forgetting and channel redundancy caused by the excessive hidden state number. To address the above challenges, we design a local block to model local connectivity and facilitate the expressive power of different channels. The local block shown in Figure 2(c), it is a simplified channel attention. Given a features $F_{in}$, the local features $F_{L}$ can be obtained by:
$F_{in}^{{}^{\prime}}$ $=F_{in}\otimes W_{p}^{(1)}(GAP(F_{in})),$ (6)
$F_{L}$ $=W_{p}^{(2)}(F_{in}^{{}^{\prime}})$
where $W_{p}^{(\cdot)}$ is the $1\times 1$ convolution, GAP is the global average pool
III-D Features Aggregation (FA)
As image details primarily consist of local features, we design a features aggregation(FA) module, as illustrated in Figure 2(b), to emphasize the importance of the local block in restoration by recalibrating the weights through the learnable factor when aggregating the two blocks for recovery. Specifically, given the global features $F_{G}$ and local features $F_{L}$, the aggregate process can be defined as:
$F_{A}=F_{G}+WF_{L}$ (7)
where $W$ denotes the learnable parameters, which are directly optimized by backpropagation and initialized as 1. It is worth mentioning that our design is extremely lightweight, as it does not introduce additional convolution layers.
IV Experiments
We first describe the experimental details of the proposed ALGNet. Then we present both qualitative and quantitative comparisons between ALGNet and other state-of-the-art methods. Following that, we conduct ablation studies to validate the effectiveness of our approach. Finally, we assess the resource efficiency of ALGNet.
IV-A Experimental Settings
In this section, we introduce the details of the used datasets, and training details.
TABLE I: Image motion deblurring results. The best and second best scores are highlighted and underlined. Our ALGNet-B and ALGNet are trained only on the GoPro dataset but achieves the highest and second-highest scores on the average of the effects on both datasets, respectively.
Methods PSNR $\uparrow$ SSIM $\uparrow$
DeblurGAN-v2 [13] 29.55 0.934
SPAIR [39] 32.06 0.953
MIMO-UNet++ [5] 32.45 0.957
MPRNet [6] 32.66 0.959
MPRNet-local [6] 33.31 0.964
HINet [7] 32.71 0.959
HINet-local [7] 33.08 0.962
Uformer [18] 32.97 0.967
MSFS-Net [40] 32.73 0.959
MSFS-Net-local [40] 33.46 0.964
NAFNet-32 [4] 32.83 0.960
NAFNet-64 [4] 33.62 0.967
Restormer [16] 32.92 0.961
Restormer-local [16] 33.57 0.966
Stripformer [17] 33.08 0.962
IRNeXt [3] 33.16 0.962
SFNet [25] 33.27 0.963
FSNet [26] 33.29 0.963
ALGNet(Ours) 33.49 0.964
ALGNet-B(Ours) 34.05 0.969
TABLE II: Image motion deblurring results.
Methods PSNR $\uparrow$ SSIM $\uparrow$
DeblurGAN [11] $24.51$ $0.871$
DeblurGAN-v2 [13] 26.61 0.875
DBGAN [12] 28.94 0.915
MT-RNN [41] 29.15 0.918
DMPHN [42] 29.09 0.924
SPAIR [39] 30.29 0.931
MIMO-UNet++ [5] 29.99 0.930
MPRNet [6] 30.96 0.939
MPRNet-local [6] 31.19 0.945
Uformer [18] 30.83 0.952
MSFS-Net [40] 31.05 0.941
MSFS-Net-local [40] 31.30 0.943
Restormer [16] 31.22 0.942
Restormer-local [16] 31.49 0.945
Stripformer [17] 31.03 0.940
SFNet [25] 31.10 0.941
FSNet [26] 31.05 0.941
ALGNet(Ours) 31.68 0.957
IV-A1 Datasets
In this section, we will introduce the datasets used and provide details regarding the training configurations.
Image Motion Deblurring. Following recent methods [26, 6], we train ALGNet using the GoPro dataset [20], which includes 2,103 image pairs for training and 1,111 pairs for evaluation. To assess the generalizability of our approach, we directly apply the GoPro-trained model to the test images of the HIDE [43] and RealBlur [44] datasets. The HIDE dataset contains 2,025 images that collected for human-aware motion deblurring. Both the GoPro and HIDE datasets are synthetically generated, but the RealBlur dataset comprises image pairs captured under real-world conditions. This dataset includes two subsets: RealBlur-J, formed with camera JPEG outputs, and RealBlur-R, generated offline through processes like white balance adjustment, demosaicking, and denoising applied to RAW images.
Single-Image Defocus Deblurring. To evaluate the effectiveness of our method, we adopt the DPDD dataset [45], following the methodology of recent approaches [26, 16]. This dataset comprises images from 500 indoor/outdoor scenes captured using a DSLR camera. Each scene consists of three defocused input images and a corresponding all-in-focus ground-truth image, labeled as the right view, left view, center view, and the all-in-focus ground truth. The DPDD dataset is partitioned into training, validation, and testing sets, comprising 350, 74, and 76 scenes, respectively. ALGNet is trained using the center view images as input, with loss values computed between outputs and corresponding ground-truth images.
IV-A2 Training details
For different tasks, we train separate models, and unless specified otherwise, the following parameters are employed. The models are trained using the Adam optimizer [46] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is set to $5\times 10^{-4}$ and gradually reduced to $1\times 10^{-7}$ using the cosine annealing strategy [47]. The batch size is chosen as $32$, and patches of size $256\times 256$ are extracted from training images. Data augmentation involves horizontal and vertical flips. We scale the network width by setting the number of channels to 32 and 64 for ALGNet and ALGNet-B, respectively.
IV-B Experimental Results
IV-B1 Image Motion Deblurring
The quantitative results on the GoPro and HIDE dataset are shown in Table I II. Our ALGNet-B shows 0.43 dB performance improvement over NAFNet-64 [4] on GoPro [20]. Moreover, as shown in Figure 1, through the scaling up of the model size, our ALGNet achieves even better performance, highlighting the scalability of ALGNet. Noted that, even though our network is trained solely on the GoPro [20] dataset, it still achieves a substantial gain of 0.19 dB PSNR over Restormer-Local [16] on the HIDE [43] dataset. This demonstrates its impressive generalization capability. Figure. 4 shows our model recovered more visually pleasant results. We also evaluate our ALGNet on the real-world images of RealBlur dataset [44] and achieve the best performance as shown in Table III.
Figure 3: Image motion deblurring comparisons on the GoPro dataset [20]. Compared to the state-of-the-art methods, our ALGNet excels in restoring sharper and perceptually faithful images.
Figure 4: Image motion deblurring comparisons on the RealBlur dataset [44].
TABLE III: Quantitative real-world deblurring results.
RealBlur-R RealBlur-J
Methods PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
DeblurGAN-v2 [13] 36.44 0.935 29.69 0.870
MPRNet [6] 39.31 0.972 31.76 0.922
Stripformer [17] 39.84 0.975 32.48 0.929
MRLRFNet [19] 40.92 0.975 33.19 0.936
FFTformer [14] 40.11 0.973 32.62 0.932
ALGNet(ours) 41.21 0.977 33.38 0.937
TABLE IV: Quantitative comparisons with previous leading single-image defocus deblurring methods on the DPDD testset [45] (containing 37 indoor and 39 outdoor scenes).
Indoor Scenes Outdoor Scenes Combined
Methods PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ LPIPS $\downarrow$
EBDB [48] 25.77 0.772 0.040 0.297 21.25 0.599 0.058 0.373 23.45 0.683 0.049 0.336
DMENet [49] 25.50 0.788 0.038 0.298 21.43 0.644 0.063 0.397 23.41 0.714 0.051 0.349
JNB [50] 26.73 0.828 0.031 0.273 21.10 0.608 0.064 0.355 23.84 0.715 0.048 0.315
DPDNet [45] 26.54 0.816 0.031 0.239 22.25 0.682 0.056 0.313 24.34 0.747 0.044 0.277
KPAC [51] 27.97 0.852 0.026 0.182 22.62 0.701 0.053 0.269 25.22 0.774 0.040 0.227
IFAN [52] 28.11 0.861 0.026 0.179 22.76 0.720 0.052 0.254 25.37 0.789 0.039 0.217
Restormer [16] 28.87 0.882 0.025 0.145 23.24 0.743 0.050 0.209 25.98 0.811 0.038 0.178
IRNeXt [3] 29.22 0.879 0.024 0.167 23.53 0.752 0.049 0.244 26.30 0.814 0.037 0.206
SFNet [25] 29.16 0.878 0.023 0.168 23.45 0.747 0.049 0.244 26.23 0.811 0.037 0.207
FSNet [26] 29.14 0.878 0.024 0.166 23.45 0.747 0.050 0.246 26.22 0.811 0.037 0.207
ALGNet(Ours) 29.37 0.898 0.023 0.147 23.68 0.755 0.048 0.223 26.45 0.821 0.036 0.186
IV-B2 Single-Image Defocus Deblurring
Table IV presents a numerical comparison of defocus deblurring methods on the DPDD dataset [45]. ALGNet outperforms other state-of-the-art methods across most metrics. Notably, in the outdoor scenes category, ALGNet exhibits a 0.15 dB improvement over the leading method IRNeXt [3]. The visual results in Figure 5 illustrate that our method recovers more details and visually aligns more closely with the ground truth compared to other algorithms.
Figure 5: Single image defocus deblurring comparisons on the DDPD dataset [45]. Compared to the state-of-the-art methods, our ALGNet effectively removes blur while preserving the fine image details.
IV-C Ablation Studies
Here we present ablation experiments to verify the effectiveness and scalability of our method. Evaluation is performed on the GoPro dataset [20], and the results are shown in Table. V. The baseline is NAFNet [4]. We perform the break-down ablation by applying the proposed modules to the baseline successively, we can make the following observations: a) The Local block and Global block do not lead to significant performance gains when used alone. This is because our baseline has been able to capture the local information well, so the local block cannot improve the representation ability of the model too much. However, when we replace the module in baseline with our proposed global block, due to the problems of pixel forgetting and channel redundancy, it cannot improve the performance significantly. b) When we use both Global Block and local Block, we can see a big improvement in performance, up 0.52dB compared to baseline. c) The FA makes a gain of 0.14 dB for our model. e) When all three of them are used together, the lifting capacity of each module is higher.
In addition, to examine the advantage of our FA design, we compare our FA with other method, such as sum and concatenation. As shown in Table VI, our FA achieves better results, which indicates that our FA is able to emphasize the importance of the local block in restoration, enabling our model to select the most detail information for recovery. It is worth mentioning that our design is extremely lightweight, as it does not introduce additional convolution layers.
TABLE V: Ablation study on individual components of the proposed ALGNet.
Global Block Local Block FA PSNR
[…]
TABLE VI: The impact of skip connections on the overall performance.
Modules Sum Concatenation FA
PSNR 33.35 33.37 33.49
FLOPs(G) 17 22 17
IV-D Resource Efficient
We examine the model complexity of the proposed approach and the state-of-the-art methods in terms of model running time, and FLOPs. Table VII and Figure 1 demonstrate that our ALGNet model attains state-of-the-art performance while concurrently reducing computational costs. Specifically, on the GoPro dataset [20], we achieve a 0.2 dB improvement over the previous best approach, FSNet [26], while using with up to 84.7% of cost reduction and achieving nearly 1.5$\times$ faster inference. This highlights the efficiency of our method, showcasing superior performance along with resource effectiveness.
TABLE VII: The evaluation of model computational complexity on the GoPro dataset [20]. The FLOPs are evaluated on image patches with the size of 256×256 pixels. The running time is evaluated on images with the size of 1280 × 720 pixels.
Method Time(s) FLOPs(G) PSNR SSIM
DBGAN [12] 1.447 759 31.10 0.942
MPRNet [6] 1.148 777 32.66 0.959
Restormer [16] 1.218 140 32.92 0.961
Stripformer [17] 1.054 170 33.08 0.962
IRNeXt [3] 0.255 114 33.16 0.962
SFNet [25] 0.408 125 33.27 0.963
FSNet [26] 0.362 111 33.29 0.963
ALGNet(Ours) 0.237 17 33.49 0.964
V Conclusion
In this paper,we propose an efficient image deblurring network that leverages selective structured state spaces model to aggregate enriched and accurate features. We design an aggregate local and global block (ALGBlock) consisting of both global and local components. The global block utilizes selective structured state spaces model to capture long-range dependency information. And the local block models local connectivity using simplified channel attention and address the issues of local pixel forgetting and channel redundancy in the state space equation. In addition, we present an aggregate methods that emphasize the importance of the local block in restoration. Experimental results demonstrate that the proposed method outperforms state-of-the-art approaches on widely used benchmarks.